# Backstepping for Partial Differential Equations

### v2

@@ 2.9.3 Complex-valued equations
While most of the early backstepping designs focused on real-valued PDEs, the methodology readily extends to complex-valued equations, encompassing models such as the Schrodinger and Ginzburg-Landau equations. These equations often arise in quantum mechanics, nonlinear optics, and fluid dynamics, and their control presents unique challenges due to the complex nature of their states and dynamics.
@@
In [186], the problem of boundary control for the linearized Schrodinger equation was tackled with backstepping. The key idea was to formally view the Schrodinger equation as a complex-valued heat equation and apply the backstepping method developed for parabolic PDEs. This resulted in a full-state feedback controller that achieved an arbitrary decay rate. The control gains were expressed in closed form, involving Kelvin functions, highlighting the explicit nature of the backstepping design.
As an illustrative example, consider the linearized Schrodinger equation:
@@
which can be written in term of Kelvin functions (see [186]) and then used as usual to obtain the control law; exponential stability of the closed-loop system is as before obtained from the stability of the target system and the invertibility of the transformation.
@@ 3 Interconnected systems
for all . If some velocities are equal, then the velocity matrix presents isotachic blocks (these are blocks of states sharing the same transport speed). This specific case is exposed in Section 3.1.3. The functions and are continuously differentiable. The in-domain coupling term is defined as , where , , and are continuous functions. Adjusting the exponential change of coordinates presented in [95], it is possible to consider that the matrices and do not have diagonal terms as there always exist canonical coordinates which allow transforming the system into such a canonical form. Most contributions in the literature consider the velocity matrix and the coupling matrix as constant to simplify the exposition, although the different computations and results can be easily extended to spatially varying matrices. We adopt this choice in the rest of this section. Finally, the boundary couplings and are constant and are respectively called distal reflection (reflection at the unactuated boundary) and proximal reflection (reflection at the actuated boundary). The matrices and are constant. Depending on the configurations considered in the different subsections, some coupling terms will be set to zero. We will also specify the control input and the available measurement.
@@
where are identity operators of adequate size (that can reasonably be extended to bounded and boundedly invertible operators), is a bounded operator, is a matrix and is a bounded operator (in most cases, a family of matrices parametrized by the spatial variable). Finally, the operator is chosen as a bounded and boundedly invertible integral operator. It is usually a Volterra integral operator of the second kind (as in (8)). However, as it will appear throughout this section, several configurations require adding affine terms (with a diagonal structure) or even considering Fredholm integral operators. In this latter case, the invertibility of the transformation is not granted anymore and should be carefully proved. Due to its block-triangular structure and to the invertibility of the blocks on the diagonal, the transformation (52) is boundedly invertible. Consequently, as explained in Section 2.6, the original system and the target system share equivalent stability properties.
@@ 3.1 Coupled Hyperbolic systems
We assume that the control law acts at the boundary, i.e. we have . Given the potentially destabilizing in-domain coupling terms in (51), a logical step in designing the backstepping target system is to neutralize these terms. More precisely, in [95], the target system is considered as a copy of the initial dynamics (51) but with and . This target system is finite-time stable. Indeed, applying the method of characteristics, we have that for all ,
@@
As shown in [95], the original system can be mapped to the desired target system using the backstepping transformation (52) (without the and states and where is a second kind Volterra integral operator). The kernel satisfies a set of hyperbolic PDEs that can be rewritten as integral equations using the method of characteristics. Then, applying the method of successive approximations, it is possible to show the existence of a unique fixed point and, consequently, a unique continuous solution to these kernel equations [119, Theorem 3.2].
@@
A collocated state-observer is proposed in [277] using the measurement . This observer is defined as a copy of the system plus injection of output error tuned using the bakcstepping approach to guarantee the exponential stability of the error system (see Section 2.7). This observer is combined with the state feedback controller (54) in [277] to obtain an output-feedback controller that stabilizes the original system (51) in finite time .
@@ 3.1.2 Homodirectional systems of equations and heterodirectional systems
Once the problem of stabilizing two coupled scalar hyperbolic equations had been solved, subsequent studies have concentrated on non-scalar heterodirectional hyperbolic systems with rightward propagating states and leftward propagating states. Such systems are referred to as heterodirectional systems. They correspond to equation (51) without ODEs. Note that or heterodirectional systems are qualified as homedirectional systems since all transport velocities have the same signs, meaning all PDEs convect in the same direction. Due to the finite length of the spatial domain, homodirectional systems are inherently stable, but the coupling between states can lead to undesirable transient behaviors, making trajectory planning problems non-trivial.
@@
An alternative target system (and therefore stabilizing control law) is proposed in [88]. It reads as
@@ 3.2 Hyperbolic systems and Integral Delay Equations
Hyperbolic systems of linear balance laws exhibit notable connections with a specific class of time-delay systems. Initially established by D’Alembert’s formula [98], which transforms wave equations into difference equations, this connection extends through the method of characteristics, allowing to represent as difference equations a class of linear conservation laws. In [239], the existence of a mapping between the solutions of first-order hyperbolic PDEs and zero-order neutral systems is proved using spectral methods. This mapping is shown to be unique and is explicitly constructed in [162] for a single hyperbolic equation with a reaction term. Interestingly, the proposed methodology emphasizes the simplicity of stability analysis as techniques initially developed for time-delay systems [150, 218] can then be applied for the analysis of hyperbolic PDEs. We show below how the backstepping approach has been used to generalize the results of [162] to heterodirectional systems. Such connections between hyperbolic systems and time-delay systems can be crucial to design stabilizing backstepping controllers for complex interconnected systems or to analyze the robustness properties of the backstepping controllers.
@@ 3.2.1 Time-delay formulation of heterodirectional systems
where is a positive integer (that depends on and ), the delays depend on the transport velocities and , the constant matrices depend on the boundary coupling term and , while the function is piecewise continuous. The state can be seen as a function that belongs to , i.e. the Banach space of functions mapping the interval into . The associated norm is defined as We refer the reader to [150] for more mathematical details on this mathematical framework. It is shown in [33] and [21, Theorem 6.1.3] that the IDE (60) and the PDE system (51) have equivalent stability properties. More precisely, there exist and two constants and , such that for any linear bounded state-feedback law , for any
@@ 3.2.2 Delay-robustness of backstepping controllers
If the open-loop transfer function of a feedback system has an infinite number of poles in the complex right-half plane, then the introduction of arbitrarily small time delays in the loop may cause instability for any feedback control law [100, 209]. Therefore, it is first essential to analyze the open-loop behavior of the system (3.1). Based on the characteristic equation (62), it is required in [33, 30] that the open-loop PDE system (51) with and in the absence of the ODE states is exponentially stable. This assumption is a necessary condition for delay-robust stabilization. It corresponds to the exponential stability of the principal part of the IDE (60) (i.e., the equation ) [150] and avoids having an infinite number of unstable poles in open loop. For , this assumption rewrites . For larger values of and , it can be verified using sufficient conditions based on Lyapunov-Krasovskii methods [224, 137].
Even when this assumption is verified, it is shown in [33] that the backstepping controllers may not be robust to input delays as they significantly trade off delay-robustness for performance to guarantee finite-time stability. More precisely, the controllers with a structure as in (54) are composed of two parts:
@@ 3.3.1 Advanced problems for time-delay systems
The backstepping approach could then be extended to non-linear ODEs with input delays. The case of multi-input nonlinear systems with constant delays is solved in [51], while the case of time-varying input delays is considered in [46] for the the system
@@
The backstepping methodology has proven valuable in analyzing the delay-robustness properties of predictor-based controllers. For linear ODEs, conditions on the constant delay mismatch that guarantee robustness are given in [176]. Recently, these results have been extended in [174] to obtain conditions guaranteeing mean-square exponential stability in the presence of a stochastic delay. For nonlinear systems, the case of time-dependent and state-dependent delay uncertainties is tackled in [49], closed-loop robustness is shown if the magnitudes of the delay uncertainty and its partial derivatives were small enough. For unknown (but constant) time delays, adaptive backstepping controllers have been introduced in [183] when an upper bound of the delay (denoted ) is known. Such adaptive control law require knowing the distributed state , which is restrictive. An input observer is designed in [64] to estimate . However, global stability is no longer guaranteed in this configuration. This control framework has been extended to systems with distinct delays [312] or distributed delays [313]. Output-feedback controller designs are addressed in [62] using dual approaches while solving input disturbance rejection. Finally, we mention that a neural network-based estimation scheme is proposed in [74] to estimate the unknown input delay magnitude.
@@ 3.3.2 Interconnected ODE-PDE systems
As we have seen above, it has been possible to design backstepping stabilizing controllers dynamical systems with input and measurement delays. Subsequently, these control techniques have been extended to design observers, controllers, or parameter estimation methods for a general class of interconnected ODE-PDE systems, for which the PDE subsystem is not a simple transport equation. Consider the system (51) in the absence of the -ODE state (i.e. ). Although different actuator configurations can be considered (as either the ODE state or the PDE state can be actuated), most contributions in the literature focus on the case of an actuated PDE. Therefore, we have . To deal with the joined presence of the ODE and the heterodirectional PDE system, the backstepping transformations found in the literature have the structure (52) (without the term and with ). In [119], such a backstepping transformation is used to map the system (51) to the exponentially stable target system
where is strictly upper-triangular. Provided that is Hurwitz, the target system (69) is exponentially stable. This target system can be seen as the natural extension of (55). Due to the presence of the ODE term in the backstepping transformation, the kernel equations rewrite as a coupled PDE-ODE system, whose well-posedness is shown in [119, Theorem 3.2]. Compared to (54), the stabilizing control input includes an additional term of the form .
@@
The main advantage of this approach is that only the conventional kernel equations have to be solved, whereas the decoupling kernel equations take a simple form. This two-step approach has been adjusted in [116] to design a state-observer. Adjustments of this multi-step approach are proposed in [142] for scalar hyperbolic systems ( coupled with ODE systems in normal forms. The proposed backstepping design exploits the strict feedback form of the system to successively recursively stabilize it. Extensions to the case of a non-linear ODE are proposed in [156]. As the proposed controller requires prediction of the ODE and PDE states, appropriate predictors are designed as the unique solution of general nonlinear Volterra integro-differential equations.
@@
Systems with state-dependent or time-dependent input delays have been addressed for instance in [124], [123], and [125] developing predictor-feedback control designs with applications to 3D printing processes, where material transport in screw extruders creates complex input delay dynamics. Finally, we emphasize that specific attention has been recently paid to wave equations coupled with ODEs [237, 112, 310].
@@ 3.3.3 Interconnected ODE-PDE-ODE systems
The first application of backstepping to the stabilization of interconnected ODE-PDE-ODE systems is found in [12]. In this contribution, the two ODEs are scalar, and the PDE system corresponds to a scalar transport equation from the -ODE to the -ODE. In other words, this contribution shows how to incorporate first-order actuator and sensor dynamics into the controller and observer designs for the scalar 1-D linear hyperbolic PDE system considered in [193]. The proposed state-feedback controller uses the transformation (56) (without the ODE terms) and does not guarantee the exponential stability of the ODE state . The case of a scalar heterodirectional PDE system () is considered in [120] for a scalar actuated ODE (i.e., ). Under appropriate stabilizability conditions for the ODE systems, the original interconnected system (51) is mapped to a target system with a similar structure but with and where and are replaced by Hurwitz matrices. The matrix is also replaced by a matrix chosen to guarantee the closed-loop exponential stability. Rather than canceling the entire proximal wave reflection in the control law, it is assigned the dynamics of a high-pass filter. Therefore, the resulting target system dynamics take the form of a cascade of the distal ODE into a first-order neutral system, which, when stable, is shown to be -stable in the sense [96]. As explained in Section 3.2.2, such a filtering approach robustifies the controller. A state-observer relying on collocated measurements is also designed in [120] using a dual approach.
@@
The limitations induced by this relative degree one condition are partially mitigated in [61] for a scalar PDE system, where the -ODE is presumed to be minimum phase for the influencing output of the PDE. This methodology was broadened in [283] to include a state observer, thereby allowing the design of an output-feedback controller. Nonetheless, the PDE subsystem is still assumed to be scalar in this contribution. The extension to the stabilization of an interconnected ODE- heterodirectional PDE-ODE system is proposed in [30], while a state observer can be found in [27]. Using the backstepping transformation (52) (with ), the system (51) is mapped to a target system that can, in turn, be rewritten as an IDE system (coupled with ODEs) with equivalent stability properties [21, Lemma 6.1.4].
@@
The first condition corresponds to the (delay-) robustness assumption stated in [33, Theorem 4] and is necessary to guarantee the delay-robustness of the control input (see Section 3.2.1). The second assumption is not very conservative since it is impossible to stabilize the ODE subsystems independently of the PDE or interconnection structure without the stabilizability of and at the very least. The last assumption implies that the matrices and are not identically zero and that does not have any zeros in that is common to all its components (and therefore admits a right inverse). Although conservative, this last assumption is less restrictive than other conditions existing in the literature and previously mentioned. The robustness of the closed-loop system is guaranteed in [30] using an appropriate filtering procedure (explained in Section 3.2.1).

### v3

@@ 2.9.3 Complex-valued equations
While most of the early backstepping designs focused on real-valued PDEs, the methodology readily extends to complex-valued equations, encompassing models such as the Schrödinger and Ginzburg-Landau equations. These equations often arise in quantum mechanics, nonlinear optics, and fluid dynamics, and their control presents unique challenges due to the complex nature of their states and dynamics.
@@
In [187], the problem of boundary control for the linearized Schrödinger equation was tackled with backstepping. The key idea was to formally view the Schrödinger equation as a complex-valued heat equation and apply the backstepping method developed for parabolic PDEs. This resulted in a full-state feedback controller that achieved an arbitrary decay rate. The control gains were expressed in closed form, involving Kelvin functions, highlighting the explicit nature of the backstepping design.
As an illustrative example, consider the linearized Schrödinger equation:
@@
which can be written in terms of Kelvin functions (see [187]) and then used as usual to obtain the control law; exponential stability of the closed-loop system is as before obtained from the stability of the target system and the invertibility of the transformation.
@@ 3 Interconnected systems
for all . If some velocities are equal, then the velocity matrix presents isotachic blocks (these are blocks of states sharing the same transport speed). This specific case is exposed in Section 3.1.3. The functions and are continuously differentiable. The in-domain coupling term is defined as , where , , and are continuous functions. Applying the exponential change of coordinates presented in [96], it is possible to consider that the matrices and do not have diagonal terms as there always exist canonical coordinates which allow transforming the system into such a canonical form. Most contributions in the literature consider the velocity matrix and the coupling matrix as constant to simplify the exposition, although the different computations and results can be easily extended to spatially varying matrices. We adopt this choice in the rest of this section. Finally, the boundary couplings and are constant and are respectively called distal reflection (reflection at the unactuated boundary) and proximal reflection (reflection at the actuated boundary). The matrices and are constant. Depending on the configurations considered in the different subsections, some coupling terms will be set to zero. We will also specify the control input and the available measurement.
@@
where are identity operators of adequate size (that can reasonably be extended to bounded and boundedly invertible operators), is a bounded operator, is a matrix and is a bounded operator (in most cases, a family of matrices parametrized by the spatial variable). Finally, the operator is chosen as a bounded and boundedly invertible integral operator. It is usually a Volterra integral operator of the second kind (as in (8)). However, as it will appear throughout this section, several configurations require adding affine terms (with a diagonal structure) or even considering Fredholm integral operators. In this latter case, the invertibility of the transformation is not granted anymore and should be carefully proved. Due to its block-triangular structure and the invertibility of the blocks on the diagonal, the transformation (52) is boundedly invertible. Consequently, as explained in Section 2.6, the original system and the target system share equivalent stability properties.
@@ 3.1 Coupled Hyperbolic systems
We assume that the control law acts at the boundary, i.e. we have . Given the potentially destabilizing coupling terms in (51), a logical step in designing the backstepping target system is to neutralize these terms. More precisely, in [96], the target system is considered as a copy of the initial dynamics (51) but with and . This target system is finite-time stable. Indeed, applying the method of characteristics, we have that for all ,
@@
As shown in [96], the original system can be mapped to the desired target system using the backstepping transformation (52) (without the and states and where is a Volterra integral operator of the second kind). The kernel satisfies a set of hyperbolic PDEs that can be rewritten as integral equations using the method of characteristics. Then, applying the method of successive approximations, it is possible to show the existence of a unique fixed point and, consequently, a unique continuous solution to these kernel equations [96, Theorem A.1]. Higher regularity of the solutions can also be shown under adequate regularity of the coefficients appearing in the kernel PDEs [96, Theorem A.2].
@@
A collocated state-observer is proposed in [282] using the measurement . This observer is defined as a copy of the system plus injection of output error tuned using the backstepping approach to guarantee the exponential stability of the error system (see Section 2.7). This observer is combined with the state feedback controller (54) in [282] to obtain an output-feedback controller that stabilizes the original system (51) in finite time .
@@ 3.1.2 Homodirectional systems of equations and heterodirectional systems
Once the problem of stabilizing two coupled scalar hyperbolic equations had been solved, subsequent studies have concentrated on non-scalar heterodirectional hyperbolic systems with rightward propagating states and leftward propagating states. Such systems are referred to as heterodirectional systems. They correspond to equation (51) without ODEs. Note that or heterodirectional systems are qualified as homodirectional systems since all transport velocities have the same signs, meaning all PDEs convect in the same direction. Due to the finite length of the spatial domain, homodirectional systems are inherently stable, but the coupling between states can lead to undesirable transient behaviors, making trajectory planning problems non-trivial.
@@
An alternative target system (and therefore stabilizing control law) is proposed in [89]. It is given by
@@ 3.2 Hyperbolic systems and Integral Delay Equations
Hyperbolic systems of linear balance laws exhibit notable connections with a specific class of time-delay systems. Initially established by D’Alembert’s formula [99], which transforms wave equations into difference equations, this connection extends through the method of characteristics, allowing to represent as difference equations a class of linear conservation laws. In [241], the existence of a mapping between the solutions of first-order hyperbolic PDEs and zero-order neutral systems is proved using spectral methods. This mapping is shown to be unique and is explicitly constructed in [163] for a single hyperbolic equation with a reaction term. Interestingly, the proposed methodology emphasizes the simplicity of stability analysis, as techniques initially developed for time-delay systems [151, 219] can then be applied for the analysis of hyperbolic PDEs. We show below how the backstepping approach has been used to generalize the results of [163] to heterodirectional systems. Such connections between hyperbolic systems and time-delay systems can be crucial to designing stabilizing backstepping controllers for complex interconnected systems or to analyzing the robustness properties of the backstepping controllers.
@@ 3.2.1 Time-delay formulation of heterodirectional systems
where is a positive integer (that depends on and ), the delays depend on the transport velocities and , the constant matrices depend on the boundary coupling term and , while the function is piecewise continuous. The state can be seen as a function that belongs to , i.e. the Banach space of functions mapping the interval into . The associated norm is defined as We refer the reader to [151] for more details on this mathematical framework. It is shown in [33] and [21, Theorem 6.1.3] that the IDE (60) and the PDE system (51) have equivalent stability properties. More precisely, there exist and two constants and , such that for any linear bounded state-feedback law , for any
@@ 3.2.2 Delay-robustness of backstepping controllers
If the open-loop transfer function of a feedback system has an infinite number of poles in the complex right-half plane, then the introduction of arbitrarily small time delays in the loop may cause instability for any feedback control law [101, 210]. Therefore, it is first essential to analyze the open-loop behavior of the system (3.1). Based on the characteristic equation (62), it is required in [33, 30] that the open-loop PDE system (51) with and in the absence of the ODE states is exponentially stable. This assumption is a necessary condition for delay-robust stabilization. It corresponds to the exponential stability of the principal part of the IDE (60) (i.e., the equation ) [151] and avoids having an infinite number of unstable poles in open loop. For , this assumption is equivalent to . For larger values of and , it can be verified using sufficient conditions based on Lyapunov-Krasovskii methods [225, 138].
Even when this assumption is verified, it is shown in [33] that the backstepping controllers may not be robust to input delays as they significantly trade-off delay-robustness for performance to guarantee finite-time stability. More precisely, the controllers with a structure as in (54) are composed of two parts:
@@ 3.3.1 Advanced problems for time-delay systems
The backstepping approach could then be extended to non-linear ODEs with input delays. The case of multi-input nonlinear systems with constant delays is solved in [51], while the case of time-varying input delays is considered in [46] for the system
@@
The backstepping methodology has proven valuable in analyzing the delay-robustness properties of predictor-based controllers. For linear ODEs, conditions on the constant delay mismatch that guarantee robustness are given in [177]. Recently, these results have been extended in [175] to obtain conditions guaranteeing mean-square exponential stability in the presence of a stochastic delay. For nonlinear systems, the case of time-dependent and state-dependent delay uncertainties is tackled in [49], and closed-loop robustness is shown if the magnitudes of the delay uncertainty and its partial derivatives are small enough. For unknown (but constant) time delays, adaptive backstepping controllers have been introduced in [184] when an upper bound of the delay (denoted ) is known. Such adaptive control law require knowing the distributed state , which is restrictive. An input observer is designed in [65] to estimate . However, global stability is no longer guaranteed in this configuration. This control framework has been extended to systems with distinct delays [317] or distributed delays [318]. Output-feedback controller designs are addressed in [63] using dual approaches while solving input disturbance rejection. Finally, we mention that a neural network-based estimation scheme is proposed in [75] to estimate the unknown input delay magnitude.
@@ 3.3.2 Interconnected ODE-PDE systems
As we have seen above, it has been possible to design backstepping stabilizing controllers for dynamical systems with input and measurement delays. Subsequently, these control techniques have been extended to design observers, controllers, or parameter estimation methods for a general class of interconnected ODE-PDE systems, for which the PDE subsystem is not a simple transport equation. Consider the system (51) in the absence of the -ODE state (i.e. ). Although different actuator configurations can be considered (as either the ODE state or the PDE state can be actuated), most contributions in the literature focus on the case of an actuated PDE. Therefore, we have . To deal with the joined presence of the ODE and the heterodirectional PDE system, the backstepping transformations found in the literature have the structure (52) (without the term and the corresponding terms in the transformation). In [120], such a backstepping transformation is used to map the system (51) to the exponentially stable target system
where is strictly upper-triangular. Provided that is Hurwitz, the target system (69) is exponentially stable. This target system can be seen as the natural extension of (55). Due to the presence of the ODE term in the backstepping transformation, the kernel equations can be rewritten as a coupled PDE-ODE system, whose well-posedness is shown in [120, Theorem 3.2]. Compared to (54), the stabilizing control input includes an additional term of the form .
@@
The main advantage of this approach is that only the conventional kernel equations have to be solved, whereas the decoupling kernel equations take a simple form. This two-step approach has been adjusted in [117] to design a state-observer. Adjustments of this multi-step approach are proposed in [143] for scalar hyperbolic systems ( coupled with ODE systems in normal forms. The proposed backstepping design exploits the strict feedback form of the system to recursively stabilize it. Extensions to the case of a non-linear ODE are proposed in [157]. As the proposed controller requires prediction of the ODE and PDE states, appropriate predictors are designed as the unique solution of general nonlinear Volterra integro-differential equations.
@@
Systems with state-dependent or time-dependent input delays have been addressed, for instance, in [125], [124], and [126] developing predictor-feedback control designs with applications to 3D printing processes, where material transport in screw extruders creates complex input delay dynamics. Finally, we emphasize that specific attention has been recently paid to wave equations coupled with ODEs [239, 113, 315].
@@ 3.3.3 Interconnected ODE-PDE-ODE systems
The first application of backstepping to the stabilization of interconnected ODE-PDE-ODE systems is found in [61], for two scalar heterodirectional transport equations () with potentially unstable ODE “actuator” and “load” dynamics, where the “actuator” ODE is required to have an invertible input matrix, and the “load” ODE is required to be stabilizable. This result obtains a target system where in-domain couplings in the PDEs are removed, as well as the proximal reflection coefficient on the PDEs (i.e., on the side of the actuated ODE) and the coupling from the PDEs to the actuated ODE. Furthermore, in the target system, the two ODEs are stabilized.
Another early contribution can be found in [12], limited to two scalar ODEs, and the PDE system corresponds to a scalar transport equation from the -ODE to the -ODE. In other words, this contribution shows how to incorporate first-order actuator and sensor dynamics into the controller and observer designs for the scalar 1-D linear hyperbolic PDE system considered in [194]. The proposed state-feedback controller uses the transformation (56) (without the ODE terms) and does not guarantee the exponential stability of the ODE state . The case of a scalar heterodirectional PDE system () is considered in [121] for a scalar actuated ODE (i.e., ). Under appropriate stabilizability conditions for the ODE systems, the original interconnected system (51) is mapped to a target system with a similar structure but with and where and are replaced by Hurwitz matrices. The matrix is also replaced by a matrix chosen to guarantee the closed-loop exponential stability. Rather than canceling the entire proximal wave reflection in the control law, it is assigned the dynamics of a high-pass filter. Therefore, the resulting target system dynamics take the form of a cascade of the distal ODE into a first-order neutral system, which, when stable, is shown to be -stable in the sense [97]. As explained in Section 3.2.2, such a filtering approach robustifies the controller. A state-observer relying on collocated measurements is also designed in [121] using a dual approach.
@@
The limitations induced by this relative degree one condition are partially mitigated in [62] for a scalar PDE system, where the -ODE is presumed to be minimum phase for the influencing output of the PDE. This article was also the first to include a strictly proper control for ODE-PDE-ODE systems, by adding an adequately chosen low-pass filter in the control design. As mentioned in previous sections, this can have a large impact on the resulting robustness of the control laws for PDE systems. This methodology was broadened in [287] to include a state observer, thereby allowing the design of an output-feedback controller. Nonetheless, the PDE subsystem is still assumed to be scalar in this contribution. The extension to the stabilization of an interconnected ODE- heterodirectional PDE-ODE system is proposed in [30], while a state observer can be found in [27]. Using the backstepping transformation (52) (with ), the system (51) is mapped to a target system that can, in turn, be rewritten as an IDE system (coupled with ODEs) with equivalent stability properties [21, Lemma 6.1.4].
@@
The first condition corresponds to the (delay-) robustness assumption stated in [33, Theorem 4] and is necessary to guarantee the delay-robustness of the control input (see Section 3.2.1). The second assumption is not very conservative since it is impossible to stabilize the ODE subsystems, independently of the PDE or interconnection structure, without the stabilizability of and at the very least. The last assumption implies that the matrices and are not identically zero and that does not have any zeros in that are common to all its components (and therefore admits a stable right inverse). Although conservative, this last assumption is less restrictive than other conditions existing in the literature and previously mentioned. The robustness of the closed-loop system is guaranteed in [30] using an appropriate filtering procedure (explained in Section 3.2.1).